OpenStudy (jymoore24):

WILL FAN AND MEDAL!!!!!! A metal worker has a metal alloy that is 20% copper and another alloy that is 60% copper. How many kilograms of each alloy should the metalworker combine to create 80 kg of a 52% copper alloy?

OpenStudy (xguardians):

Take a look at this: https://www.algebra.com/algebra/homework/word/mixtures/Mixture_Word_Problems.faq.question.900309.html It may help you find your solution.

OpenStudy (aaronandyson):

A = 20% copper alloy. B = 60% copper alloy. a + b = 80 a = (80-b) So get 52% copper. We multiply .52 by 80 that gives 41.6 So we may frame this equation: .20a + .60b = .52(80) .20a + .60b = 41.6 Now we substitute a = (80-b) in this equation. .20(80-b) + .60b = 41.6 16 - .20b + .60b = 41.6 .40b = 41.6 - 16 b = 25.6/.4 b = 64 kg of 60% copper : Now can you calculate a? To make it easier for you I'll do the first step i.e frame the frame equation. Remember this: a + b = 80 b = 64 kg of 60% copper(Calculated). a + 64 = 80 a = 80-64 a= 16 kg of 20% copper.